# Boundary‑Informed Method of Lines for Physics‑Informed Neural Networks

### v1

@@ Abstract
[…]
Figure 2.1: Side-by-side comparison of the standard PINN and the MOL-PINN architectures.
@@ 3 Results
Figure 3.1: *
Table 3.1: Hyper-parameters and hardware. Model Epochs LR Parameter 8 000 $1.0\!\times\!10^{-3}$ Model Optim GPU Parameter Adam RTX 4070 Epochs, Learning Rate, Optimizer, and GPU
Figure 3.2: *
Figure 3.3: *
Table 3.2: MOL-PINN MSE results & comparison. PINN 21i 50i 100i Burgers $7.15\times 10^{-3}$ $2.40\times 10^{-1}$ $1.27\times 10^{-2}$ $2.79\times 10^{-3}$ NS–TG $5.87\times 10^{-1}$ $2.23\times 10^{-2}$ $2.17\times 10^{-2}$ N/A NS–ABC $1.169\times 10^{0}$ $7.74\times 10^{-1}$ N/A N/A Emphasizing the affect of variable $i$ MOL trajectories
Figure 3.4: *
@@
Table 3.3: Regional MSE for 3D/4D Navier–Stokes
$v$ $u$ $p$ (TG) / $w$ (ABC) PINN TG $1.00\times 10^{-2}$ $6.58\times 10^{-3}$ $1.75\times 10^{0}$ MOL-PINN TG $2.61\times 10^{-2}$ $2.35\times 10^{-2}$ $1.56\times 10^{-2}$ PINN ABC $54\times 10^{0} 1.1$$15\times 10^{0}$ $1.01\times 10^{0}$ MOL-PINN ABC $1.29\times 10^{0}$ $6.78\times 10^{-1}$ $3.54\times 10^{-1}$
Figure 3.5: 3D Navier-Stokes Error Heatmap (50i v-region)
@@ 3.2 Time-based Efficiency
Figure 3.2: Burgers MSE Heatmap (100i)
Table 3.4: Time per epoch for tested simulations
PINN 21i 50i 100i Darcy 0.012s 0.046s 0.142s N/A Burgers 0.0142s 0.017s 0.046s 1.015s Navier-Stokes 3D 0.013s 0.278s 0.367s N/A Navier-Stokes 4D 0.031s 2.25s N/A N/A
Figure 3.3: N-S Output Heatmap (50i v-region)

### v2

# Maximilian Cederholm

@@ 1 Introduction
Figure 1: Visual comparison between PINN and MOL-PINN architectures.
3 Hardware and Hyper-parameters
All the relevant training hardware and hyperparameters are summarized in table 1.
Table 1: Hyper-parameters and hardware.
Parameter Value
Epochs 8,000
Learning Rate (LR) $1.0\times 10^{-3}$
Optimizer Adam
GPU RTX 4070
@@ 4 Results
Table 2: MOL-PINN MSE results & comparison.
PINN 21i 50i 100i
Burgers $7.15\times 10^{-3}$ $2.40\times 10^{-1}$ $1.27\times 10^{-2}$ $2.79\times 10^{-3}$
NS–TG $5.87\times 10^{-1}$ $2.23\times 10^{-2}$ $2.17\times 10^{-2}$ N/A
NS–ABC $1.169\times 10^{0}$ $7.74\times 10^{-1}$ N/A N/A
Table 3: *
Emphasizing the effect of variable $i$ MOL trajectories.
Figure 2: Image 1 caption
Figure 3: Image 2 caption
Figure 4: Image 3 caption
@@ 4.1 N-S Results
Table 4: Regional MSE for 3D/4D Navier–Stokes
$v$ $u$ $p$ (TG) / $w$ (ABC)
PINN TG $1.00\times 10^{-2}$ $6.58\times 10^{-3}$ $1.75\times 10^{0}$
MOL-PINN TG $2.61\times 10^{-2}$ $2.35\times 10^{-2}$ $1.56\times 10^{-2}$
PINN ABC $54\times 10^{0} 1.1$$15\times 10^{0}$ $1.01\times 10^{0}$
MOL-PINN ABC $1.29\times 10^{0}$ $6.78\times 10^{-1}$ $3.54\times 10^{-1}$
@@ 4.2 Time-based Efficiency
Table 5: Time per epoch for tested simulations
PINN 21i 50i 100i
Darcy 0.012s 0.046s 0.142s N/A
Burgers 0.0142s 0.017s 0.046s 1.015s
Navier-Stokes 3D 0.013s 0.278s 0.367s N/A
Navier-Stokes 4D 0.031s 2.25s N/A N/A